Problem 3.27 Your answer is incorrect. Try again At t = 484 s after midnight, a spacecraft of mass 1200 kg is located at position <3 x 105, 4 × 105-3 × 105> m, and at that time an asteroid whose mass is 7 x 1015 kg is located at position <9 x 105, -3 x 105, -17 x 105>m. There are no other objects nearby. (a) Calculate the (vector) force acting on the spacecraft. F net0.036 8.29e5 1.66e4 > N (b) At t = 484 s the spacecrafts momentum was pi, and at the later time ț = 491 s its momentum was pf. Calculate the (vector) change of momentum p-P pf-pi = || 0.0000546 58.03e5 11.62e4 > kg.m/s
Answer #1

Distance between the asteroid and spacecraft which will be given as -

| \vec{r}| = \sqrt{}(x2 - x1)2 + (y2 - y1)2 + (z2 - z1)2

| \vec{r}| = \sqrt{}[(9 x 105 m) - (3 x 105 m)]2 + [(-3 x 105 m) - (4 x 105 m)]2 + [(-17 x 105 m) - (-3 x 105 m)]2

| \vec{r}| = \sqrt{}2.81 x 1012 m2

| \vec{r}| = 1.67 x 106 m

The unit vector can be calculated by -

\hat{r} = \vec{r} / | \vec{r}|

\hat{r} = (x2 - x1) \hat{i} + (y2 - y1) \hat{j} + (z2 - z1) \hat{k} / \sqrt{}(x2 - x1)2 + (y2 - y1)2 + (z2 - z1)2

\hat{r} = [(9 x 105 m) - (3 x 105 m)] \hat{i} + [(-3 x 105 m) - (4 x 105 m)] \hat{j}+ [(-17 x 105 m) - (-3 x 105 m)] \hat{k} / (1.67 x 106 m)

\hat{r} = (6 x 105 m) \hat{i} + (-7 x 105 m) \hat{j} + (-14 x 105 m) \hat{k} / (1.67 x 106 m)

\hat{r} = (0.359) \hat{i} + (-0.419) \hat{j} + (-0.838) \hat{k}

(a) The force acting on the spacecraft which will be given by -

\vec{F}net = {G M m / | \vec{r}|2 } \hat{r}

where, M = mass of a spacecraft = 1200 kg

m = mass of an asteroid = 7 x 1015 kg

G = gravitational constant = 6.67 x 10-11 Nm2/kg2

then, we get

\vec{F}net = { [(6.67 x 10-11 Nm2/kg2) (1200 kg) (7 x 1015 kg)] / (2.81 x 1012 m2) } [(0.359) \hat{i} + (-0.419) \hat{j} + (-0.838) \hat{k}]

\vec{F}net = (1.99 x 10-4 N) [(0.359) \hat{i} + (-0.419) \hat{j} + (-0.838) \hat{k}]

\vec{F}net = (7.14 x 10-5 N) \hat{i} + (-8.33 x 10-5 N) \hat{j} + (-1.66 x 10-4 N) \hat{k}

Then, we have

\vec{F}net = < 7.14 x 10-5 , -8.33 x 10-5 , -1.66 x 10-4 > N